1. Sketch the 3 trigonometric functions of sin θ, cos θ and tan θ over one complete cycle, where 00 ≤ θ ≤ 360 

 

Complete the sketches on separate graphs for the three trigonometric functions. 


1
Expert's answer
2021-02-24T06:34:07-0500

"0\leq\theta\leq360\degree"


"y=\sin \theta, Period=360\degree"
"y=\cos \theta, Period=360\degree"





"y=\tan \theta, Period=180\degree,"

"\text{Vertical\ asymptotes}\ \theta=90\degree+180\degree n, n\in\Z"